Investor X has 7,000 to spend on the following three projects. Use the AW method to determine which of these independent investments are financially acceptable at 6% per year compounded monthly interest rate. Lifelong probl so Use CC=AW/i to get present worth.

Option

Installed Cost

Return per month

A

4500

220

B

3000

200

C

2200

140
Answer #1

Monthly worth
A:
=220-4500*6%/12=197.5

B:
=200-3000*6%/12=185

C:
=140-2200*6%/12=129

Accept B and C